OpenStudy (anonymous):

By using triple integral, find the volume of the solid bounded by the cylinder y^2+4z^2=16 and planes x=0 and x+y =4

OpenStudy (anonymous):

It would be first convenient to convert the first equation to polar coordinates. $$y^2 + 4z^2 = 16 \\ \frac 1 4 y^2 + z^2 = 4$$ Try it.